OpenStudy (anonymous):

finding general solution and finding theta (1)

OpenStudy (anonymous):

\[d \theta / dt = \sqrt{\theta +1}\]

OpenStudy (anonymous):

Multiply both sides by dt then integral both sides then use u-substitution on the right side of the equation.

OpenStudy (anonymous):

2/3(\theta +1)^2/3 +C\]

OpenStudy (sirm3d):

\[\Large \frac{d\theta}{\sqrt{\theta+1}}=dt\] use the substitution \[\Large u=\theta+1,\quad du=d\theta\]